Two numbers are consecutive odd integers. The sum of the smaller and 4 times the larger is 63. Find the two numbers.
Answer by ccs2011(207)

n + 4(n+2) = 63
n + 4n + 8 = 63
5n + 8 = 63
5n = 55
n = 11
The 2 numbers are 11 and 13